Cavalier accounting firm can produce, on average,  2 tax returns per day.

The operation has a design capacity of 16 tax returns per day and an effective capacity of 14 tax returns per day. (Round answer to 1 decimal place.

     

  Utilization %
  Efficiency %

     

b.

Browns Carpet Cleaning can clean an average of 4 carpets each day if the design capacity is 8 carpets per day and the effective capacity is 7 carpets per day. (Round your answer to 1 decimal place)

      

  Utilization %
  Efficiency %
Answer #1

a)

Utilization = Actual Output / Design Capacity = 2/16 = 12.5%
Efficiency = Actual output / Effective capacity = 2/14 = 14.3%

b)

Utilization = Actual Output / Design Capacity = 4/8 = 50.0%
Efficiency = Actual output / Effective capacity = 4/7 = 57.1%